What is the arc length of an arc with central angle 262 degrees and radius 9cm? Express your answer in pi and round the final answer to the nearest hundredth.
Answer by stanbon(75887)

What is the arc length of an arc with central angle 262 degrees and radius 9cm? Express your answer in pi and round the final answer to the nearest hundredth.
---
circumference of the circle = 2(pi)r = 2(pi)9 = 56.55 cm
----
Set up a proportion:
x/262 = 56.55/360
---
x = 262[56.55/360]
---
x = 41.15 (the arc length)
======================
Cheers,
stan H.